# Emergence of topological states in relaxation dynamics of interacting bosons

### v1

@@ Abstract
Topological concepts have been employed to understand the ground states of many strongly correlated systems, but it is still quite unclear if and how topology manifests itself in the relaxation dynamics. Here we uncover emergent topological phenomena in the time evolution of far-from-equilibrium one-dimensional interacting bosons. Beginning with simple product states, the system evolves into long-time stationary states with high energy that are nonthermal for a wide range of parameters, and they exhibit nonlocal string correlation that is characteristic of the symmetry-protected topological ground state of the Hamiltonian. In contrast, no topological feature is found in the stationary state as long as the system thermalizes. This difference is further corroborated by the distinct behaviour of quantum entanglement and edge states of the system. Our theoretical prediction can be examined by current experimental techniques and paves the way for a more comprehensive understanding of topological phases in nonequilibrium settings.
@@
In non-relativistic quantum mechanics, the time evolution of an isolated interacting system is deterministic as dictated by its Hamiltonian [1, 2, 3, 4]. It is expected that relaxation dynamics of a nonequilibrium initial state results in a steady state whose local observables are described by the generalized Gibbs ensemble [5, 6, 7, 8] or Gibbs ensemble [9, 10, 11, 12]. Experimental investigations of relaxation dynamics are challenging because a sufficiently large number of quantum objects should be isolated from the environment and precisely manipulated for a considerable amount of time. In this regard, ultracold atomic gases provide a versatile platform for exploring nonequilibrium quantum dynamics, where the system can be driven far from equilibrium by sudden changes in the Hamiltonian [13]. To date, most studies along this direction employed many-body systems that are trivial from the topological perspective [14, 15, 16, 17, 18, 19, 20, 21].
@@
It is by now well established that the properties of some many-body states should be described using topological concepts. The most common setting is to explore topology in the ground states of many-body systems that are well separated from the excited states [22, 23, 24, 25, 26]. Remarkable experimental progresses on ground-state topology of interacting systems have been reported in several quantum simulators [27, 28, 29, 30]. However, far-from-equilibrium dynamics and long-time stationary state of interacting topological systems are largely unexplored [31, 32, 33, 34, 35]. For an isolated system, its total energy remains constant during time evolution [36]. Consequently, the time-evolved stationary state is a high-lying excited state that has substantial overlap with an appreciable number of eigenstates (see Fig. 1a), and the physics is expected to be quite different from the ground state. It raises the question as to what extent topological fingerprints of the Hamiltonian can be revealed during time evolution, and whether an isolated system initially prepared in a nonequilibrium state can reach a new equilibrium with interesting topological phenomena. These issues are also intimately connected with ongoing experiments using ultracold atoms.
@@
Figure 1: Emergent topology in the relaxation dynamics of an interacting system. a, The most commonly studied scenario of topological phases is based on the properties of the many-body ground state that are separated from the excited states by a gap. In contrast, it is not transparent if and how the high-lying excited states that form a continuum can be defined as topological. b, Schematic of the one-dimensional superlattice Bose-Hubbard model and emergence of topological states that are far from the ground state in its relaxation dynamics. The system has alternating hopping $J_{1,2}$ and onsite repulsion $U$ and is initialized in certain simple product states (upper panel). Depending on the values of $J_{1,2}$ and $U$, the long-time stationary state may be topologically nontrivial with long-range string correlation or trivial without such correlation (lower panel).
@@
In this paper, we provide compelling evidence for the emergence of topological states in the far-from-equilibrium relaxation dynamics of interacting bosons in a finite one-dimensional superlattice. As illustrated in Fig. 1b, the lattice potential is modulated such that the tunneling between two neighboring sites have different ratios depending on the height of the potential barrier. For our purpose, the system can be described by a superlattice Bose-Hubbard model with alternating hopping terms and onsite interactions. This is a bosonic analogue of the Su-Schrieffer-Heeger model [37], whose many-body ground state has been studied in several previous works [38, 39]. Using the staggered-immersion cooling method [40, 41], the system can be initially prepared in certain product states that are far from equilibrium and topologically trivial. For a wide range of parameters, the system relaxes and reaches nonthermal stationary states that have nontrivial topological properties reminiscent of the many-body ground state, as revealed by nonlocal string correlation, entanglement, and edge state. An experimental protocol is proposed to verify our theoretical results using currently available techniques.
Results
@@
Model. We study nonequilibrium dynamics of interacting bosons in a superlattice shown in Fig. 1b. Each unit cell contains two lattice sites with equal potential depths such that the system possesses bond-inversion symmetry. In the tight-binding limit, it is described by the superlattice Bose-Hubbard Hamiltonian [38]
@@
with $\delta\hat{n}_{i}=\bar{n}-\hat{n}_{i}$ ($\bar{n}$ is the average particle density) [44, 45, 46]. The expectation value with respect to the ground state remains finite and approaches a nonzero constant as $|i-j|\rightarrow\infty$ in the SPT phase [47, 48].
Time-evolved stationary state. In the limits of $J_{1,2}/U\rightarrow 0$ and $J_{1}/J_{2}\rightarrow 0$ or $J_{2}/J_{1}\rightarrow 0$, the many-body Hamiltonian in Eq. (1) is integrable and time evolution is trivial. Instead, we focus on the nonequilibrium dynamics in the regime where $J_{1,2}/U$ and $J_{1}/J_{2}$ are finite, where one may naively expect that the system thermalizes in the long-time limit. The system is initialized in trivial product states $\left|\psi_{0}\right\rangle=\left|a_{1}a_{2}\cdots\right\rangle$, where the particle number on each site $a_{i}=0,1$ ($\sum_{i}a_{i}=N$). Time evolution is computed using exact diagonalization (ED) for small system sizes [49] and the time dependent variational principle (TDVP) in the framework of matrix product states (MPS) for large system sizes [50, 51, 52]. We choose open boundary condition for the system and focus on the half-filled cases with the number of bosons being $N=L/2$.
Figure 2: Relaxation dynamics in topologically nontrivial and trivial cases. a, Schematics of the topologically trivial product states $\left|1010\cdots\right\rangle$ (green) and $\left|1100\cdots\right\rangle$ (orange) that are used as initial states in our simulations. b-e, Numerical results for the nontrivial case with $J_{1}/J_{2}=0.2$ and $U/J_{2}=8$. b, The string correlation $O_{2,L-1}(t)$ of the time-evolved stationary state attains an appreciable value, which is distinct from the thermal averaged value $O_{2,L-1}^{\rm thermal}\approx 0$ at the mean energy of the initial state (represented by the dashed line). c, Time evolution of the second-order Renyi entanglement entropy $S_{\mathcal{A}}(t)$ at the center of the $L=32$ system. d, The Renyi entropy at $t=100/J_{2}$ and the thermal entropy given by the canonical ensemble for the $L=32$ system. The discrepancy suggests that the system has not thermalized. e, $O_{2,L-1}(t)$ has quite weak dependence on $L$ so its nonzero value should not be finite-size effects. f-i, Numerical results for the trivial case with $J_{1}/J_{2}=0.8$ and $U/J_{2}=6$. Each panel presents the same quantity as its counterpart in panels b-e. The time-evolved stationary state has no long-range string correlation, which is consistent with the value $O_{2,L-1}^{\rm thermal}\approx 0$ for a thermalized system. The time and subsystem size dependence of the Renyi entropy also suggest that thermalization occurs. Finite-size effects are also excluded by studying multiple $L$. Numerical simulations are performed for system size up to $L=60$ and bond dimension up to $3000$ [53].
We first consider a quench from the initial state $\left|\psi_{0}\right\rangle=\left|1010\cdots\right\rangle$ or $\left|1100\cdots\right\rangle$ (see Fig. 2a) using a Hamiltonian whose ground state is topological ($J_{1}/J_{2}=0.2$, $U/J_{2}=8$). It is found that the many-body system relaxes and reaches a long-time stationary state that is far from the ground state. Surprisingly, the time-evolved high-lying state has long-range string order. As shown in Fig. 2b, the correlation between the second and $(L-1)$-th sites $O_{2,L-1}(t)=\left\langle\psi(t)\right|\hat{O}_{2,L-1}\left|\psi(t)\right\rangle$ oscillates at the beginning but quickly attains a stable value with negligible fluctuations. The variation of $O_{i,j}(t)$ with $|i-j|$ is also consistent with the existence of long-range order [53]. This behaviour is observed for all system sizes that we have studied (see Fig. 2e), so it should be a genuine phenomenon but not a finite-size effect. For the TDVP calculations, multiple bond dimensions have been tested such that the maximal one captures the entanglement of the time-evolved state [53]. For the small $L=12$ system, long-range string order is observed for time up to $t=1000/J_{2}$ using ED [53].
The existence of long-range string correlation is associated with the absence of thermalization. At the energy corresponding to the initial state $\left|\psi_{0}\right\rangle$, thermal average of the string correlation is $O_{2,L-1}^{\rm thermal}\approx 0$. If the system reaches thermal equilibrium, $O_{2,L-1}(t)$ should approach zero in the long-time limit, but we clearly obtain nonzero values depending on the initial states. Thermalization can also be probed using the second-order Renyi entanglement entropy $S_{\mathcal{A}}(t)=-\text{log}\left[\text{Tr}\left(\rho^{2}_{\mathcal{A}}(t)\right)\right]$, where $\mathcal{A}$ is a subsystem with $l_{\mathcal{A}}$ sites and $\rho_{\mathcal{A}}(t)$ is the associated reduced density matrix. The half-system entropy increases with time and begins to saturate to a constant at the late time $t=100/J_{2}$ (see Fig. 2c). The saturated entropy is noticeably smaller than the thermal entropy given by canonical ensemble as shown in Fig. 2d, which suggests that the time-evolved stationary state is nonthermal [54].
Next we study the topologically trivial case with $J_{1}/J_{2}=0.8$ and $U/J_{2}=6$. In contrast to previous results, the system does thermalize and memory about the initial states is lost when it reaches equilibrium. In Fig. 2f and i, the string correlation $O_{2,L-1}(t)$ quickly decays to nearly zero values for both initial states. The dynamics of $S_{\mathcal{A}}(t)$ for $l_{\mathcal{A}}=L/2$ is presented in Fig. 2g. It increases with time linearly for a certain interval and eventually reaches a maximum value that is constrained by the MPS bond dimension. The subsystem size dependence of $S_{\mathcal{A}}(t=100/J_{2})$ is displayed in Fig. 2h. It grows linearly with $l_{\mathcal{A}}$ and agrees with the thermal entropy given by the canonical ensemble [15]. These features clearly distinguish this case from the previous one.
It is intriguing that the nonequilibrium phase transition and thermalization is absent for hardcore interaction $U=\infty$ but can only be observed in the soft-core case [53]. To further understand this phenomenon, perturbation theory is employed to construct an effective model at half-filling for the $J_{1,2}/U\ll 1$ regime in which terms up to second order of $J_{1,2}/U$ are kept. Time evolution using the effective model and the original model in Eq. (1) agrees with each other remarkably well in the soft-core cases [53].
In addition to the string order parameter, we have also studied other quantities of the stationary state to substantiate its topological nature. It is found that the parity and density-wave orders vanish in the bulk so the long-range string correlation is not caused by some trivial states [53]. For two adjacent lattice sites $i$ and $i+1$, their reduced density matrix $\rho_{i,i+1}(t)$ can be used to define the two-site entanglement entropy $S_{i}(t)=-\text{Tr}\left[\rho_{i,i+1}(t)\text{ln}\rho_{i,i+1}(t)\right]$ [55]. In the nontrivial case, $S_{i}(t)$ oscillates between large and small values with a period of two sites for time up to $t=100/J_{2}$ (see Fig. 3a,c). When the system is trivial, $S_{i}(t)$ reaches its maximum value quickly and is basically homogeneous in real space (see Fig. 3b,d). These features are analogous to those of the many-body ground state [44, 45, 46] so the time-evolved state shares similar topological features as the ground state. The dynamics of entanglement spectrum has also been investigated and multiple level crossings are observed [53], which is similar to previous results on free fermions [56].
Figure 3: Dynamics of the two-site entropy. a, For the nontrivial case with $J_{1}/J_{2}=0.2$ and $U/J_{2}=8$, the two-site entropy oscillates between large and small values on odd and even lattice sites in the bulk of the system. b, For the trivial case with $J_{1}/J_{2}=0.8$ and $U/J_{2}=6$, the two-site entropy is almost homogeneous as one would expect for a thermal state. The line cut at $t=100/J_{2}$ in panels a and b are shown respectively in panels c and d. The same quantity for the nontrivial many-body ground state is shown in panel c for comparison. The oscillations in the ground state and the nontrivial time-evolved stationary state are similar. All panels are for the $L=36$ system.
Experimental protocol. Our proposal can be readily realized using ultracold bosons trapped in a one-dimensional spin-dependent optical superlattice [57]. The inhomogeneity induced by the Gaussian beams can be compensated by a tailored light profile to generate a flat box potential with hard walls [58, 59]. The superlattice should be adjusted to create balanced double wells so the system is described by Eq. (1) with alternating tunneling strengths $J_{1}$ and $J_{2}$ controlled by the superlattice depth. It is essential to prepare the system in simple product states at half-filling with the odd (even) sites singly occupied (empty) (see Fig. 2a), which has already been achieved in previous experiments using the staggered-immersion cooling method [40, 41].
Figure 4: Nonequilibrium phase diagram of the superlattice Bose-Hubbard model. a, Color map of the string correlation $O_{2,L-1}(t=100/J_{2})$ in the $L=36$ system for a wide range of $J_{1}/J_{2}$ and $U/J_{2}$. The point A (B) indicates the topologically nontrivial (trivial) case that has been studied in detail. b-c, Time evolution of the imbalance $\Delta_{e}(t)$ at the edge and $\Delta_{b}(t)$ in the bulk. The two quantities exhibit very different behavior for the A point but are similar for the B point. d, The long-time limit of $O_{2,L-1}(t)$ as a function of the hopping strengths for two different interactions.
To further assess the experimental feasibility, we have explored the nonequilibirum phase diagram for a wide range of parameters. By inspecting the string order parameter at $t=100/J_{2}$ (typical timescale realized experimentally), it is found that topologically nontrivial stationary states appear in a large portion of the phase diagram, as shown in Fig. 4a,d. This suggests that nontrivial relaxation dynamics should be quite accessible. In view of the single-site resolution provided by quantum gas microscope [60], we expect that nonlocal string correlation [28] and entanglement entropy [15] can be measured in future experiments.
In addition, time evolution of the edge states could also reveal useful information, as experiments on superconducting circuits and trapped ions have already demonstrated [31, 35]. It should be fruitful to measure the imbalance $\Delta_{e}(t)\equiv n_{\text{1}}(t)-n_{\text{L}}(t)\neq 0$ associated with the two edges and the averaged imbalance $\Delta_{b}(t)\equiv\frac{2}{L-2}\sum^{L-1}_{i=2}[n_{\text{i,odd}}(t)-n_{\text{i,even}}(t)]=0$ in the bulk, where $n_{i}(t)=\langle\psi(t)|{\hat{n}}_{i}|\psi(t)\rangle$. If the system reaches topologically nontrivial stationary state, $\Delta_{e}(t)$ has an appreciable value throughout the time evolution but $\Delta_{b}(t)$ decays to zero after a short period (see red lines in Fig. 4b,c). This can be taken as evidence for topologically protected edge states in the far-from-equilibrium dynamics. In contrast, both quantities rapidly decay to negligible values in the trivial case (see blue lines in Fig. 4b,c).
Method
@@
ED and TDVP are employed to study the relaxation dynamics of our system. The cases with $L\leq 12$ are studied using the ED programs in the Quspin package [49]. The two-site and one-site hybrid TDVP is applied to the cases with $L>12$ [50]. This method is favourable because it preserves unitarity for a considerable amount of time using moderate bond dimensions [51, 52]. In the two-site steps, we choose truncation error $\varepsilon=10^{-10}$ and time step $J\delta t=0.05$. Each lattice site is allowed to host at most four bosons. The canonical ensemble properties are computed using the matrix product purification method [61, 62], where a finite-temperature state is obtained by imaginary time evolution from an initial state with infinite temperature. Both TDVP and finite temperature simulations are performed using the ITensor library [63].
Discussion
@@
We have uncovered surprising emergence of topological states in the far-from-equilibrium long-time evolution of strongly interacting systems. This study goes beyond previous works that mainly studied topological systems consist of non-interacting particles [64, 65] or short-time evolution of interacting systems [66, 67, 68], and is related to ongoing experiments on ultracold atoms. Topologically distinct behaviour in the relaxation dynamics are observed as a trivial initial state can evolve to a stationary state that exhibits topological properties in certain parameter regimes. While product states were used as initial states here, other high-lying initial states may also lead to intricate topological physics. It is also possible to induce nonequilibrium topological critical phenomena and study their universality classes and scaling laws. Time evolution and nonequilibrium topology of interacting systems in higher dimensions should also be very interesting. Finally, whether and how nonlocal observables thermalize after a sufficiently long time in large systems calls for more investigations [7, 69]. These questions are very difficult to address from either the experimental or theoretical side and beyond the scope of this work.
Acknowledgements

### v2

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

@@ Abstract
Topological concepts have been employed to understand a wide range of phenomena in physics. While a plethora of ground states are known to have topological underpinning, it is still quite unclear if and how topology manifests itself in the relaxation dynamics of strongly correlated many-body systems. In this paper, we study time evolution of interacting bosons in a finite one-dimensional superlattice and uncover surprising emergent topological phenomena in the long-time stationary states. Beginning with simple product states, the system evolves into stationary states with high energy whose string correlation and entanglement are investigated. Based on extensive numerical simulations and effective-model analysis, it is demonstrated that the stationary states are nonthermal for a wide range of parameters, and they exhibit certain features that are characteristic of the symmetry-protected topological ground state of the Hamiltonian. In contrast, no topological feature is found in the stationary state as long as the system thermalizes. This difference is further corroborated by the distinct behaviour of quantum entanglement and edge states of the system. Our theoretical prediction can be examined by current experimental techniques and paves the way for a more comprehensive understanding of topological phases in nonequilibrium settings.
@@
Introduction. In non-relativistic quantum mechanics, the time evolution of an isolated interacting system is deterministic as dictated by its Hamiltonian [1, 2, 3, 4]. It is expected that relaxation dynamics of a nonequilibrium initial state results in a steady state whose local observables are described by the generalized Gibbs ensemble [5, 6, 7, 8] or Gibbs ensemble [9, 10, 11, 12]. Experimental investigations of relaxation dynamics are challenging because a sufficiently large number of quantum objects should be isolated from the environment and precisely manipulated for a considerable amount of time. In this regard, ultracold atomic gases provide a versatile platform for exploring nonequilibrium quantum dynamics, where the system can be driven far from equilibrium by sudden changes in the Hamiltonian [13]. To date, most studies along this direction employed many-body systems that are trivial from the topological perspective [14, 15, 16, 17, 18, 19, 20, 21].
@@
It is by now well established that the properties of some many-body states should be described using topological concepts. The most common setting is to explore topology in the ground states of many-body systems that are well separated from the excited states [22, 23, 24, 25, 26]. Remarkable experimental progresses on ground-state topology of interacting systems have been reported in several quantum simulators [27, 28, 29, 30]. However, far-from-equilibrium dynamics and long-time stationary state of interacting topological systems are largely unexplored [31, 32, 33, 34, 35]. For an isolated system, its total energy remains constant during time evolution [36]. Consequently, the time-evolved stationary state is a high-lying excited state that has substantial overlap with an appreciable number of eigenstates [see Fig. 1(a)], and the physics is expected to be quite different from the ground state. It raises the question as to what extent topological fingerprints of the Hamiltonian can be revealed during time evolution, and whether an isolated system initially prepared in a nonequilibrium state can reach a new equilibrium with interesting topological phenomena. These issues are also intimately connected with ongoing experiments using ultracold atoms.
@@
Figure 1: Emergent topology in the relaxation dynamics of an interacting system. (a) The most commonly studied scenario of topological phases is based on the properties of the many-body ground state that are separated from the excited states by a gap. In contrast, it is not transparent if and how the high-lying excited states that form a continuum can be defined as topological. (b) Schematic of the one-dimensional superlattice Bose-Hubbard model and emergence of topological states that are far from the ground state in its relaxation dynamics. The system has alternating hopping $J_{1,2}$ and onsite repulsion $U$ and is initialized in certain simple product states (upper panel). Depending on the values of $J_{1,2}$ and $U$, the long-time stationary state may be topologically nontrivial with long-range string correlation or trivial without such correlation (lower panel).
@@
In this paper, we provide compelling evidence for the emergence of topological states in the far-from-equilibrium relaxation dynamics of interacting bosons in a finite one-dimensional superlattice. As illustrated in Fig. 1(b), the lattice potential is modulated such that the tunneling between two neighboring sites have different ratios depending on the height of the potential barrier. For our purpose, the system can be described by a superlattice Bose-Hubbard model with alternating hopping terms and onsite interactions. This is a bosonic analogue of the Su-Schrieffer-Heeger model [37], whose many-body ground state has been studied in several previous works [38, 39]. Using the staggered-immersion cooling method [40, 41], the system can be initially prepared in certain product states that are far from equilibrium and topologically trivial. For a wide range of parameters, the system relaxes and reaches nonthermal stationary states that have nontrivial topological properties reminiscent of the many-body ground state, as revealed by nonlocal string correlation, entanglement, and edge state. Previous works mainly focused on topological response of the noninteracting systems before relaxing back to equilibrium with the time-evolved state normally being trivial [42, 43, 44, 45, 46, 47, 48, 49, 50, 51], or short-time quench dynamics of the interacting topological systems [52, 53, 54, 55]. In contrast, we reveal an unknown interaction-induced many-body phenomenon in long-time relaxation dynamics, i.e. a topologically nontrivial stationary state emerges from a trivial product state. An experimental protocol is proposed to verify our theoretical results using currently available techniques.
@@
Model and method. We study nonequilibrium dynamics of interacting bosons in a superlattice shown in Fig. 1(b). Each unit cell contains two lattice sites with equal potential depths such that the system possesses bond-inversion symmetry. In the tight-binding limit, it is described by the superlattice Bose-Hubbard Hamiltonian [38]
@@
with $\delta\hat{n}_{i}=\bar{n}-\hat{n}_{i}$ ($\bar{n}$ is the average particle density) [58, 59, 60]. The expectation value with respect to the ground state remains finite and approaches a nonzero constant as $|i-j|\rightarrow\infty$ in the SPT phase [61, 62]. The phase boundary between topologically trivial and nontrivial phases is $J_{1}/J_{2}=1$ for the noninteracting case, but is shifted to $J_{1}/J_{2}<1$ by many-body interactions [63]. In subsequent calculations, we mainly focus on two representative cases, which are $J_{1}/J_{2}=0.2,U=8$ in the topological regime and $J_{1}/J_{2}=0.8,U=6$ in the trivial regime.
@@
To study the many-body relaxation dynamics of the interacting system, we employ both exact diagonalization (ED) and the time dependent variational principle (TDVP) in the framework of matrix product states (MPS). The cases with $L\leq 12$ are studied using the ED programs in the Quspin package [80]. The two-site and one-site hybrid TDVP is applied to the cases with $L>12$ [81]. The TDVP method is favourable because it preserves unitarity for a considerable amount of time using moderate bond dimensions [82, 83]. In the two-site steps, we choose truncation error $\varepsilon=10^{-10}$ and time step $J\delta t=0.05$. Each lattice site is allowed to host at most four bosons. The canonical ensemble properties are computed using the matrix product purification method [84, 85], where a finite-temperature state is obtained by imaginary time evolution from an initial state with infinite temperature. Both TDVP and finite temperature simulations are performed using the ITensor library [86].
Time-evolved stationary state. In the limits of $J_{1,2}/U\rightarrow 0$ and $J_{1}/J_{2}\rightarrow 0$ or $J_{2}/J_{1}\rightarrow 0$, the many-body Hamiltonian in Eq. (1) is integrable and time evolution is trivial. Instead, we focus on the nonequilibrium dynamics in the regime where $J_{1,2}/U$ and $J_{1}/J_{2}$ are finite, where one may naively expect that the system thermalizes in the long-time limit. The system is initialized in trivial product states $\left|\psi_{0}\right\rangle=\left|a_{1}a_{2}\cdots\right\rangle$, where the particle number on each site $a_{i}=0,1$ ($\sum_{i}a_{i}=N$). We choose open boundary conditions for the system and focus on the half-filled cases with the number of bosons being $N=L/2$.
Figure 2: Relaxation dynamics in topologically nontrivial and trivial cases. (a) Schematics of the topologically trivial product states $\left|1010\cdots\right\rangle$ (green) and $\left|1100\cdots\right\rangle$ (orange) that are used as initial states in our simulations. (b-e) Numerical results for the nontrivial case with $J_{1}/J_{2}=0.2$ and $U/J_{2}=8$. (b) The string correlation $O_{2,L-1}(t)$ of the time-evolved stationary state attains an appreciable value, which is distinct from the thermal averaged value $O_{2,L-1}^{\rm thermal}\approx 0$ at the mean energy of the initial state (represented by the dashed line). (c) Time evolution of the second-order Renyi entanglement entropy $S_{\mathcal{A}}(t)$ at the center of the $L=32$ system. (d) The Renyi entropy at $t=100/J_{2}$ and the thermal entropy given by the canonical ensemble for the $L=32$ system. The discrepancy suggests that the system has not thermalized. (e) $O_{2,L-1}(t)$ has quite weak dependence on $L$ so its nonzero value should not be finite-size effects. (f-i) Numerical results for the trivial case with $J_{1}/J_{2}=0.8$ and $U/J_{2}=6$. Each panel presents the same quantity as its counterpart in panels (b-e). The time-evolved stationary state has no long-range string correlation, which is consistent with the value $O_{2,L-1}^{\rm thermal}\approx 0$ for a thermalized system. The time and subsystem size dependence of the Renyi entropy also suggest that thermalization occurs. Finite-size effects are also excluded by studying multiple $L$. Numerical simulations are performed for system size up to $L=60$ and bond dimension up to $3000$ [63].
In light of the state-of-the-art experimental method for state preparation, we first consider a quench from the initial state $\left|\psi_{0}\right\rangle=\left|1010\cdots\right\rangle$ or $\left|1100\cdots\right\rangle$ [see Fig. 2(a)] using a Hamiltonian whose ground state is topological ($J_{1}/J_{2}=0.2$, $U/J_{2}=8$). It is found that the many-body system relaxes and reaches a long-time stationary state that is far from the ground state. Surprisingly, the time-evolved high-lying state has long-range string order. As shown in Fig. 2(b), the correlation between the second and $(L-1)$-th sites $O_{2,L-1}(t)=\left\langle\psi(t)\right|\hat{O}_{2,L-1}\left|\psi(t)\right\rangle$ oscillates at the beginning but quickly attains a stable value with negligible fluctuations. The variation of $O_{i,j}(t)$ with $|i-j|$ is also consistent with the existence of long-range order [63]. This behaviour is observed for all system sizes that we have studied [see Fig. 2(e)], so it should be a genuine phenomenon but not a finite-size effect. Moreover, the string order is remarkably robust when the initial state is varied [63]. For the TDVP calculations, multiple bond dimensions have been tested such that the maximal one captures the entanglement of the time-evolved state [63]. For the small $L=12$ system, long-range string order is observed for time up to $t=1000/J_{2}$ using ED [63].
The existence of long-range string correlation is associated with the absence of thermalization. At the energy corresponding to the initial state $\left|\psi_{0}\right\rangle$, thermal average of the string correlation is $O_{2,L-1}^{\rm thermal}\approx 0$. If the system reaches thermal equilibrium, $O_{2,L-1}(t)$ should approach zero in the long-time limit, but we clearly obtain nonzero values depending on the initial states. Thermalization can also be probed using the second-order Renyi entanglement entropy $S_{\mathcal{A}}(t)=-\text{log}\left[\text{Tr}\left(\rho^{2}_{\mathcal{A}}(t)\right)\right]$, where $\mathcal{A}$ is a subsystem with $l_{\mathcal{A}}$ sites and $\rho_{\mathcal{A}}(t)$ is the associated reduced density matrix. The half-system entropy increases with time and begins to saturate to a constant at the late time $t=100/J_{2}$ [see Fig. 2(c)]. The saturated entropy is noticeably smaller than the thermal entropy given by canonical ensemble as shown in Fig. 2(d), which suggests that the time-evolved stationary state is nonthermal [87].
Next we study the topologically trivial case with $J_{1}/J_{2}=0.8$ and $U/J_{2}=6$. In contrast to previous results, the system does thermalize and memory about the initial states is lost when it reaches equilibrium. In Fig. 2(f) and (i), the string correlation $O_{2,L-1}(t)$ quickly decays to nearly zero values for both initial states. The dynamics of $S_{\mathcal{A}}(t)$ for $l_{\mathcal{A}}=L/2$ is presented in Fig. 2(g). It increases with time linearly for a certain interval and eventually reaches a maximum value that is constrained by the MPS bond dimension. The subsystem size dependence of $S_{\mathcal{A}}(t=100/J_{2})$ is displayed in Fig. 2(h). It grows linearly with $l_{\mathcal{A}}$ and agrees with the thermal entropy given by the canonical ensemble [15]. These features clearly distinguish this case from the previous one.
It is intriguing that the nonequilibrium phase transition and thermalization is absent for hardcore interaction $U=\infty$ but can only be observed in the soft-core case [63]. To further understand this phenomenon, perturbation theory is employed to construct an effective model at half-filling for the $J_{1,2}/U\ll 1$ regime in which terms up to second order of $J_{1,2}/U$ are kept. Time evolution using the effective model and the original model in Eq. (1) agrees with each other remarkably well in the soft-core cases [63]. But for the hardcore case, it is instead found that both local and nonlocal observables exhibit persistent oscillations up to the time $t=100/J_{2}$ [63].
Figure 3: Dynamics of the two-site entropy. (a) For the nontrivial case with $J_{1}/J_{2}=0.2$ and $U/J_{2}=8$, the two-site entropy oscillates between large and small values on odd and even lattice sites in the bulk of the system. (b) For the trivial case with $J_{1}/J_{2}=0.8$ and $U/J_{2}=6$, the two-site entropy is almost homogeneous as one would expect for a thermal state. The line cut at $t=100/J_{2}$ in panels (a) and (b) are shown respectively in panels (c) and (d). The same quantity for the nontrivial many-body ground state is shown in panel (c) for comparison. The oscillations in the ground state and the nontrivial time-evolved stationary state are similar. All panels are for the $L=36$ system.
In addition to the string order parameter, we have also studied other quantities of the stationary state to substantiate its topological nature. It is found that the parity and density-wave orders vanish in the bulk so the long-range string correlation is not caused by some trivial states [63]. For two adjacent lattice sites $i$ and $i+1$, their reduced density matrix $\rho_{i,i+1}(t)$ can be used to define the two-site entanglement entropy $S_{i}(t)=-\text{Tr}\left[\rho_{i,i+1}(t)\text{ln}\rho_{i,i+1}(t)\right]$ [88]. In the nontrivial case, $S_{i}(t)$ oscillates between large and small values with a period of two sites for time up to $t=100/J_{2}$ [see Fig. 3(a)(c)]. When the system is trivial, $S_{i}(t)$ reaches its maximum value quickly and is basically homogeneous in real space [see Fig. 3(b)(d)]. These features are analogous to those of the many-body ground state [58, 59, 60] so the time-evolved state shares similar topological features as the ground state. The dynamics of entanglement spectrum has also been investigated and multiple level crossings are observed [63], which is similar to previous results on free fermions [47].
Experimental protocol. Our proposal can be readily realized using ultracold bosons trapped in a one-dimensional spin-dependent optical superlattice [89]. The inhomogeneity induced by the Gaussian beams can be compensated by a tailored light profile to generate a flat box potential with hard walls [90, 91]. The superlattice should be adjusted to create balanced double wells so the system is described by Eq. (1) with alternating tunneling strengths $J_{1}$ and $J_{2}$ controlled by the superlattice depth. It is essential to prepare the system in simple product state at half-filling in which every two sites connected by the $J_{2}$ hopping in the bulk has only one boson [63]. Actually, high-fidelity preparation of our initial state, whose odd (even) sites are singly occupied (empty) [see Fig. 2(a)], has already been achieved in previous experiments using the staggered-immersion cooling method [40, 41].
Figure 4: Nonequilibrium phase diagram of the superlattice Bose-Hubbard model. (a) Color map of the string correlation $O_{2,L-1}(t=100/J_{2})$ in the $L=36$ system for a wide range of $J_{1}/J_{2}$ and $U/J_{2}$. The point A (B) indicates the topologically nontrivial (trivial) case that has been studied in detail. (b)(c) Time evolution of the imbalance $\Delta_{e}(t)$ at the edge and $\Delta_{b}(t)$ in the bulk. The two quantities exhibit very different behavior for the A point but are similar for the B point. (d) The long-time limit of $O_{2,L-1}(t)$ as a function of the hopping strengths for two different interactions.
To further assess the experimental feasibility, we have explored the nonequilibirum phase diagram for a wide range of parameters. By inspecting the string order parameter at $t=100/J_{2}$ (typical timescale realized experimentally), it is found that topologically nontrivial stationary states appear in a large portion of the phase diagram, as shown in Fig. 4(a)(d). This suggests that nontrivial relaxation dynamics should be quite accessible. In view of the single-site resolution provided by quantum gas microscope [92], we expect that nonlocal string correlation [28] and entanglement entropy [15] can be measured in future experiments.
In addition, time evolution of the edge states could also reveal useful information, as experiments on superconducting circuits and trapped ions have already demonstrated [31, 35]. It should be fruitful to measure the imbalance $\Delta_{e}(t)\equiv n_{\text{1}}(t)-n_{\text{L}}(t)\neq 0$ associated with the two edges and the averaged imbalance $\Delta_{b}(t)\equiv\frac{2}{L-2}\sum^{L-1}_{i=2}[n_{\text{i,odd}}(t)-n_{\text{i,even}}(t)]=0$ in the bulk, where $n_{i}(t)=\langle\psi(t)|{\hat{n}}_{i}|\psi(t)\rangle$. If the system reaches topologically nontrivial stationary state, $\Delta_{e}(t)$ has an appreciable value throughout the time evolution but $\Delta_{b}(t)$ decays to zero after a short period [see red lines in Fig. 4(b)(c)]. This can be taken as evidence for topologically protected edge states in the far-from-equilibrium dynamics. In contrast, both quantities rapidly decay to negligible values in the trivial case [see blue lines in Fig. 4(b)(c)].
@@
Discussion. We have uncovered surprising emergence of topological states in the far-from-equilibrium long-time evolution of strongly interacting systems. This study goes beyond previous works that mainly studied topological systems consist of noninteracting particles [42, 43, 50, 51] or short-time evolution of interacting systems [52, 53, 54], and is related to ongoing experiments on ultracold atoms. Topologically distinct behaviour in the relaxation dynamics are observed as a trivial initial state can evolve to a stationary state that exhibits topological properties in certain parameter regimes. While product states were used as initial states here, other high-lying initial states may also lead to intricate topological physics. It is also possible to induce nonequilibrium topological critical phenomena and study their universality classes and scaling laws. Time evolution and nonequilibrium topology of interacting systems in higher dimensions should also be very interesting. Finally, whether and how nonlocal observables thermalize after a sufficiently long time in large systems calls for more investigations [7, 93]. These questions are very difficult to address from either the experimental or theoretical side and beyond the scope of this work.